# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

### v1

@@ 5.3.4 Initial stress regime
The orientation on the principal components of the stress tensor in undisturbed conditions is another key factor for predicting the fault stability during the reservoir management. This is analyzed in Scenarios #6. The modeling outcome revealed that a rotation of the maximum horizontal stress by 90 (scenario #6a) does not provide a significant variation with respect to the reference case. By distinction, an initial stress regime close to normally consolidated conditions (scenario #6b) significantly increases the fault reactivation likelihood, as it can be appreciated from the $t_{80}$ values in Figure 17. Although the occurrence of such a scenario is unlikely, it is important to consider its implications. With the undisturbed stress regime of scenario #5b, the limit $\chi_{max}=1$ is reached at an earlier point of the PP stage on faults F1 and F2 (l.s. 6) and F4 and F5 (l.s. 8), resulting in a larger fault area that is prone to a potential reactivation (Figure 17). The earlier activation during PP causes a reactivation of faults F1 and F2 at the end of the CGI and UGS stages as well.
Figure 17: Scenarios #6: $t_{80}$ in time on fault F2. A zoom for the UGS stage is provided on the right panel of the figure.
5.4 Ranking factors favoring fault reactivation
The objective of this section is to define criteria to rank the potential of inducing “unexpected” seismic events for the different investigated scenarios. The analysis focuses on fault F3, which is the most stressed in the realistic condition of a compartment offset, and fault F2, representing the behavior at the reservoir boundaries. The results are presented in Table 3. The ranking process takes into account the following quantities from the most to the least important one:
@@ 5.4 Ranking factors favoring fault reactivation
    the maximum value of average sliding $\delta_{avg}$;
@@
UGS $\delta_{avg}$ [m] year UGS $\delta_{avg}$ [m] year
@@
6b 0.96 0.018 6 1.00 5a 0.53 $0$ 0.53
2d 0.84 0.008 9 1.00 5b 0.53 $0$ 0.53
3b 0.81 0.010 7 1.00 2a/b 0.32 $0$ 0.46
2c 0.80 0.007 9 1.00 1 0.00 $0$ 0.00
4a 0.79 0.010 8 1.00 6a 0.00 $0$ 0.00
3a 0.78 0.008 8 1.00 6b 0.00 $0$ 0.00
1 0.78 0.007 8 1.00 3a 0.00 $0$ 0.00
5a 0.78 0.007 9 1.00 3b 0.00 $0$ 0.00
2a/b 0.77 0.007 8 1.00 3c 0.00 $0$ 0.00
6a 0.74 0.007 10 1.00 3d 0.00 $0$ 0.00
3d 0.70 0.009 8 1.00 4a 0.00 $0$ 0.00
4b 0.78 0.005 - 0.78 4b 0.00 $0$ 0.00
@@
As expected, fault activation during CGI and UGS cycles exhibits distinct critical factors for boundary and central faults. For fault F2, stability is primarily influenced by the initial stress regime of the system, geomechanical properties, like reservoir stiffness, and fault characteristics, such as cohesion, static friction angle, and the presence of slip weakening. In contrast, for fault F3 geometric parameters play a major role. The stability of fault F3 is strongly influenced by a compartment offset, non-vertical fault planes, and differential pressure changes in the two compartments.
@@ 5.5 Sensitivity analysis: Stage 2
In the second stage of the sensitivity analysis, we combine the most influential factors for yielding a potential fault reactivation, as identified in Stage 1, and test the new scenarios for the storage of different gases, namely CH4, CO2, H2, and N2. We remember that the primary objective of this analysis is to explain realistic configurations likely to be encountered in the Rotliegend formation, rather than to explore “extreme” conditions. Therefore, taking into account the most influential factors from Stage 1, we chose to focus on the combinations presented in Table 4. They combine a compartment offset, non-vertical fault planes, and variations in the fault characteristics, i.e., cohesion (C1) and static friction angle (C2). Two dedicated combinations analyze the effect of different geomechanical properties, i.e., decreasing (W) and increasing (H) of the reservoir stiffness.
@@
W $E$ decreased by 30% during injection
H $E$ increased by 30% during injection
@@
The potential risk of fault reactivation is evaluated for each scenario by the parameter $t_{80}$, whose behavior in time is shown in Figure 18. Regardless of the fluid, Figure 18 shows that the combination C2 provides the realistic conditions yielding the larger $t_{80}$ values. This scenario is characterized by a compartment offset and a reduced friction angle. Conditions close to reactivation occur during both the CGI and the UGS/UHS, or long-term storage, stages, especially if the faults already experienced a reactivation during the PP. As compared to other critical combined configurations, such as C1, scenario C2 is more critical primarily because of the reduction of the friction angle.
Figure 18: Sensitivity analysis, Stage 2: $t_{80}$ in time for fault F2 and F3 and the different fluids under investigation.
Because of its significance, the outcomes of the critical combination C2 have been separately described for CH4 and the other gases. This distinction is made due to the differences in pressure history (see Figure 6).
@@ 5.5.1 UGS critical scenario
The combination of factors in configuration C2 causes a potential reactivation mainly during primary production. The condition is particularly critical for faults F2 and F3 during the UGS phase, with both faults reaching a critical conditions not only at the end of the UGS storage phase (UGS-S) but also during withdrawal (UGS-P). Figure 19, which shows $\chi_{max}$ in time for scenario C2 as compared to the reference case (scenario #1), highlights that critical conditions for fault reactivation develop earlier during PP. $\chi_{max}$ is also greater than, or close to, 0.8 during CGI and UGS. Because of the loss of symmetry, F2 is more stressed than F1 and $\chi_{max}$ achieves values larger than 0.8 on fault F3 as well. Figure 20 provides a 3D view of $\chi$ distribution in space at a few significant loading steps.
Figure 19: Scenario C2: $\chi_{max}$ in time for all faults as compared to the reference case (scenario #1).
Figure 20: Scenario C2: $\chi$ distribution in space on the fault system at the end of PP (left), CGI (middle) and UGS (right).
The behavior of the maximum displacement ($\delta_{max}$) for all the faults is shown in Figure 21. During the CGI phase, all the faults slide, primarily during the period between loading step 11 and 12, in the opposite direction to what occurred during primary production. Additionally, faults F1, F2, and F3 slightly slide again at the end of the UGS injection phase.
Figure 21: Scenario C2: maximum sliding $\delta_{max}$ in time.
@@ 5.5.2 CO2, H2, N2 storage dynamics and fault activation
The goal of this section is to compare how the temporary or permanent storage of different gases may influence the fault stability as measured by $\chi_{max}$ in the critical scenario C2. Since the pore pressure variation during PP is the same for each fluid, $\chi$ behaves identically until l.s. 10. After that, the impact on fault stability must be considered based on whether temporary or permanent storage scenarios occur.
For the permanent storage scenarios, i.e., CO2 and N2, at the beginning of the storage $\chi_{max}$ decreases due to the unloading of the shear stress caused by reservoir expansion. However, as the storage continues, the shear stress is loaded with the opposite sign, reaching a relative maximum at the end of the injection stage. This point corresponds to the full pore pressure recovery (Figure 22). As far as it concerns the storage/production cycles, $\chi_{max}$ on F1, F2 and F4-F5 reaches its maximum at the end of the UHS injection stage, corresponding to the highest value of pressure in the reservoir. For F3, $\chi_{max}$ exhibits an additional peak at the end of the UHS withdrawal stage.
Figure 22: Scenario C2: $\chi_{max}$ in time for CO2, N2 and H2.
5.5.3 Geo-chemical effects on fault stability
The inclusion geochemical effects is also considered. Based on the literature review summarized in Section 3, we have incorporated the possibility to account for geochemical influences by introducing a hardening/weakening behavior of reservoir rocks. Specifically, during the injection the Young modulus is increased or decreased by 30%, depending on the fluid involved. For CO2, both weakening and hardening scenarios have been observed, depending on the actual rock mineralogical composition. Conversely, H2 induces a weakening effect, while N2 shows a relatively neutral geo-chemical impact on the reservoir rock.
Figure 23 shows the behavior of $\chi_{max}$ for the weakening/hardening scenarios (W/H). Fault F3 remains stable throughout the simulation because of symmetry. During CO2 injection with a hardening behavior, $\chi_{max}$ first decreases and then increases with respect to the reference case. The opposite occurs for the weakening scenario. Similar considerations hold for H2 in the weakening scenario, with an even smaller difference with respect to the reference case. Based on these results, we can preliminarily conclude that geochemical effects on the reservoir rock have a minor impact on fault stability.
Figure 23: Scenario W/H: $\chi_{max}$ in time for CO2 (upper) and H2 (lower).
@@ 6 Discussion
A few papers addressing the topics of this project have recently been published. The papers by Wassing et al. (2017) and Haug et al. (2018) are of particular interest because they focus on Rotliegend reservoirs in the Netherlands and northern Germany, providing insights into fault reactivation and fault rupture during primary production. Although the authors: i) do not consider induced seismicity during UGS stages, ii) use a simplified 2D setting, and iii) employ a different modeling approach concerning fault activation, these studies carry out a parametric analysis on the same geometric/geomechanical features investigated in this project. Therefore, it is interesting to see if our results are consistent with those presented in these publications.
@@
The findings of our work mostly align with the observations by Wassing et al. (2017) regarding fault failure mechanisms and slip initiation. Our study confirms that fault slip initiates at the top of the reservoir rocks where the initial shear capacity utilization SCU, corresponding to our safety factor $\chi$, is maximum. Both studies agree that faults without offset experience reactivation at later stages of depletion, and fault rupture does not extend upwards into the caprock. Scenarios #2c, #2d, and C2, for example, confirm that in the case of a block offset, an earlier fault activation is favored and remains bounded within the reservoir depth range. Regarding the influence of contrast in elastic properties distribution, our study partially agrees with Wassing et al. (2017). They observed a smaller peak in shear stress at the bottom of the footwall reservoir block and a localized decrease in shear stresses and SCU at the bottom of the hanging wall reservoir block. From a qualitative point of view, we agree. It must be considered that Wassing et al. (2017) assigned a uniform density and elastic properties to all rocks. Our study, in particular scenario #4a and #4b, demonstrates that the contrast in elastic properties may significantly impact on stress concentration and, therefore, fault failure.
@@
Comparing our results with the findings presented by Haug et al. (2018) regarding factors influencing fault reactivation and criticality, we agree that the initial stress regime (scenario #6a and #6b) play significant roles in fault behavior. Indeed, fault criticality largely increases as the horizontal components of the natural stress regime decrease relative to the vertical stress. Furthermore, we confirm that the stiffness contrast between the reservoir and surrounding rocks (scenario #4a and #4b) governs the distribution of production-induced loading and the degree of stress rotation, impacting fault reactivation with pressure depletion. We demonstrated that stiffness contrast between rocks is a significant factor in stress redistribution, favoring fault reactivation when the contrast increases. According to Haug et al. (2018) the depletion of thicker reservoir horizons results in a stronger fault-loading compared to the depletion of thinner reservoir horizons, as the thick reservoir undergoes a relatively larger strain for the same pore fluid pressure decrease. Although our sensitivity analysis did not directly consider reservoir thickness as a parameter, we observed consistent outcomes when the larger strain is attributed to a larger pressure decrease (scenario #5b).
@@
However, there are a few areas where our project disagrees with findings from Haug et al. (2018). While they proposed that the fault throw should be half of the reservoir thickness to obtain maximum shear stress ratio (SSR) values, we found that the most critical condition occurred when the fault offset is equal to the entire reservoir thickness. These discrepancies can be attributed to variations in the modeling setup, including the horizontal-to-vertical ratio of the natural stress components and the reservoir depth. These disparities emphasize the importance of modeling approaches and setup parameters and warrant further researches to fully understand faults behavior in different geological contexts.

### v4

@@ 3.3.4 Initial stress regime
The orientation of the principal components of the stress tensor in undisturbed conditions is another key factor for predicting fault stability during reservoir management. This is analyzed in Scenarios #6. In the reference scenario, the in-situ effective stress field is defined by M1 = $\sigma_{h}$$\sigma_{v}$ = 0.64 and M2 = $\sigma_{H}$$\sigma_{v}$ = 0.83. This configuration implies a stress hierarchy of $\sigma_{v}$ >> $\sigma_{H}$ >> $\sigma_{h}$, which corresponds to a normal fault regime, consistent with field observations reported in the Tab. 2, with $\sigma_{H}$ applied orthogonal to the F4-F5 faults and $\sigma_{h}$ applied orthogonal to the F1-F2 faults. To investigate the influence of stress orientation on fault reactivation, we simulate in Scenario #6a a 90 rotation of the principal horizontal stress directions. This test serves two main purposes: (i) assessing the sensitivity of the fault system to changes in the orientation of the stress tensor, and (ii) accounting for the uncertainty in the relative alignment between the regional stress field and the fault network, so that $\sigma_{H}$ becomes orthogonal to faults F1–F2 instead.
The modeling outcome revealed that a rotation of the maximum horizontal stress by 90, corresponding to a fault system rotated in the opposite direction with respect to the regional stress field (i.e., $\sigma_{H}$ applied orthogonal to F1-F2), does not provide a significant variation with respect to the reference case. By distinction, an initial stress regime close to normally consolidated conditions (scenario #6b) significantly increases the fault reactivation likelihood, as can be seen from the $t_{80}$ values in Figure 15. Although the occurrence of such a scenario is unlikely, it is important to consider its implications. With the undisturbed stress regime of scenario #5b, the limit $\chi_{max}=1$ is reached at an earlier point of the PP stage on F1 and F2 faults (loading step 6) and F4 and F5 (loading step 8), resulting in a larger fault area that is prone to potential reactivation (Figure 15). The earlier activation during PP causes a reactivation of F1 and F2 faults, at the end of the CGI and UGS stages as well.
Figure 15: Scenarios #6: $t_{80}$ over time on fault F2. A zoom for the UGS stage is provided on the right panel of the figure. Scenario #6 includes two in-situ stress–regime variants: (#6a) stress rotation, (#6b) modified horizontal stress ratios ($M_{1}=0.40$, $M_{2}=0.47$).
3.3.5 Ranking factors favoring fault reactivation
The objective of this section is to define criteria to rank the potential of inducing “unexpected” seismic events for the different investigated scenarios. The analysis focuses on fault F3, which is the most stressed in the realistic condition of a compartment offset, and fault F2, which represents the behavior at the reservoir boundaries.
The ranking process takes into account the following quantities
@@ 3.3.5 Ranking factors favoring fault reactivation
    the maximum value of average sliding $d_{avg}$;
@@
and follows a hierarchical comparison, giving priority to the $\chi_{max}$ indicator and using the others only when needed to discriminate between cases. The results are presented in Table 5.
@@
UGS $d_{avg}$ [m] year $P$ UGS $d_{avg}$ [m] year $P$
@@
6b 0.96 0.018 6 1.00 5a 0.53 0.000 - 0.53
2d 0.84 0.008 9 1.00 5b 0.53 0.000 - 0.53
3b 0.81 0.010 7 1.00 2a/b 0.32 0.000 - 0.46
2c 0.80 0.007 9 1.00 1 no activation due to symmetry
4a 0.79 0.010 8 1.00 6a no activation due to symmetry
3a 0.78 0.008 8 1.00 6b no activation due to symmetry
1 0.78 0.007 8 1.00 3a no activation due to symmetry
5a 0.78 0.007 9 1.00 3b no activation due to symmetry
2a/b 0.77 0.007 8 1.00 3c no activation due to symmetry
6a 0.74 0.007 10 1.00 3d no activation due to symmetry
3d 0.70 0.009 8 1.00 4a no activation due to symmetry
4b 0.78 0.005 - 0.78 4b no activation due to symmetry
@@
As expected, fault activation during CGI and UGS cycles exhibits distinct critical factors for boundary and central faults. For fault F2, stability is mainly influenced by the initial stress regime of the system, geomechanical properties, such as reservoir stiffness, and fault characteristics, such as cohesion, static friction angle, and the presence of slip weakening. In contrast, for fault F3, geometric parameters play a major role. The stability of fault F3 is strongly influenced by a compartment offset, non-vertical fault planes, and differential pressure changes in the two compartments.
@@ 3.4 Sensitivity analysis: Stage 2
In the second stage of the sensitivity analysis, we combine the most influential factors to produce a potential fault reactivation, as identified in Stage 1, and test the new scenarios for the storage of different gases, namely CH4, CO2, H2, and N2. The scenarios discussed here are summarized in Table 6. We recall that the primary objective of this analysis is to explain realistic configurations likely to be encountered in the Rotliegend formation, rather than to explore “extreme” conditions by combining the most unfavorable parameter values identified in Stage 1. Therefore, taking into account the most influential factors from Stage 1, we chose to focus on geologically plausible combinations presented in Table 6. They combine a compartment offset, non-vertical fault planes, variations in the fault characteristics, i.e., cohesion and static friction angle, and reservoir stiffness. Two dedicated combinations analyze the effect of different geomechanical properties, i.e., decreasing (W) and increasing (H) of reservoir stiffness most likely associated to chemo-mechanical effects after the PP phase.
@@
C3 $c=0$ MPa, $\delta=+65^{\circ}$, $o=100$ m, $E=20$ GPa
C4 $\varphi_{s}=20^{\circ}$, $\delta=+65^{\circ}$, $o=100$ m, $E=20$ GPa
W $E$ decreased by 30% after PP
H $E$ increased by 30% after PP
@@
The potential hazard of fault reactivation is evaluated for each combination by the parameter $t_{80}$, whose behavior in time is shown in Figure 16. Regardless of the fluid, Figure 16 shows that the combination C2 provides the conditions yielding the larger $t_{80}$ values. Therefore, all subsequent considerations will focus on this specific scenario. Scenario C2 is characterized by a compartment offset and a reduced friction angle. Conditions close to reactivation occur during both CGI and UGS/UHS, or long-term storage, stages. Notice that $t_{80}$ after the end of PP is larger if the faults already experienced a significant reactivation (i.e., with a large $t_{80}$) during PP. As compared to other critical combined configurations, scenario C2 is more critical primarily because of the reduction of the friction angle.
Figure 16: Sensitivity analysis, Stage 2: $t_{80}$ over time for fault F2 and F3 and the different fluids under investigation. The legend refers to the combinations of settings addressed in Stage 2 of the sensitivity analysis and reported in Table 6
Due to its significance, the outcomes of the critical combination C2 have been separately described for CH4 and the other gases. This distinction is made based on the differences in pressure history (see Figure 6).
@@ 3.4.1 UGS critical scenario
The combination of factors in configuration C2 causes a potential reactivation, mainly during PP. During UGS, faults F2 and F3 reach near-critical values of $\chi_{max}$ at the end of both the withdrawal and storage phases. Figure 17, which shows $\chi_{max}$ in time for scenario C2 as compared to the reference case (scenario #1), highlights that critical conditions for fault reactivation develop earlier during PP. $\chi_{max}$ is also greater than, or close to, 0.8 during CGI and UGS. Because of the loss of symmetry, F2 is more stressed than F1 and $\chi_{max}$ achieves values larger than 0.8 on fault F3 as well. Figure 18 provides a 3D view of $\chi$ distribution in space at a few significant loading steps while Figure 19 shows the time behaviour of the maximum sliding $d_{m}ax$.
Figure 17: Scenario C2, UGS: $\chi_{max}$ over time for all faults compared to the reference case (scenario #1). Scenario C2 corresponds to the low-friction configuration with $\varphi_{s}=20^{\circ}$, fault dip $\delta=+65^{\circ}$, and Block 2 offset of 100 m.
Figure 18: Scenario C2: $\chi$ distribution in space on the fault system at the end of PP (a), CGI (b) and UGS (c).
The behavior of $d_{max}$ for all the faults is shown in Figure 19. During the CGI phase, all the faults slide, primarily during the period between loading step 11 and 12, in the opposite direction to what occurred during PP. Additionally, faults F1, F2, and F3 slightly slide again at the end of the UGS injection phase.
Figure 19: Scenario C2, UGS: maximum sliding $d_{max}$ over time for all the faults.
@@ 3.4.2 CO2, H2, N2 storage dynamics and fault activation
The goal of this section is to compare how the temporary or permanent storage of different gases may influence the fault stability, as measured by $\chi_{max}$ in the critical scenario C2. Since the pore pressure variation during PP is the same for each fluid, $\chi$ has the same behavior until loading step 10. After that, the impact on fault stability must be considered based on whether temporary or permanent storage scenarios occur.
For the permanent storage scenarios, i.e., CO2 and N2, at the beginning of the storage $\chi_{max}$ decreases due to the unloading of the shear stress caused by reservoir expansion. However, as the storage continues, the shear stress is loaded with the opposite sign, reaching a relative maximum at the end of the injection stage. This point corresponds to the full pore pressure recovery (Figure 20). As far as it concerns the injection/withdrawal cycles for H2, $\chi_{max}$ on F1, F2 and F4-F5 reaches its maximum at the end of the injection stage, corresponding to the highest value of pressure in the reservoir. For F3, $\chi_{max}$ exhibits an additional peak at the end of the UHS withdrawal stage.
Figure 20: Scenario C2: $\chi_{max}$ over time for CO2, N2 and H2 after the end of PP.
3.4.3 Geochemical effects on fault stability
Due to the uncertainty observed in the literature review (Section 1.2), geochemical effects are considered in a simplified, indirect manner. We account for potential fluid-dependent influences on reservoir rock stiffness by modifying the Young modulus. Specifically, after PP the Young modulus is increased or decreased by 30%, depending on the fluid involved. For CO2, both weakening and hardening scenarios have been observed, depending on the actual rock mineralogical composition. Conversely, H2 might induce a weakening effect, while N2 shows a relatively neutral geochemical impact on the reservoir rock.
Figure 21 shows the behavior of $\chi_{max}$ for the weakening/hardening scenarios (W/H). Fault F3 remains stable throughout the simulation due to its vertical dip and the absence of offset, which limit the development of shear stress on the fault plane. During CO2 injection, with a hardening behavior, $\chi_{max}$ first decreases and then increases with respect to the reference case. The opposite occurs for the weakening scenario. Similar considerations hold for H2 in the weakening scenario, with an even smaller difference with respect to the reference case. Based on these results, we can preliminarily conclude that geochemical effects on the reservoir rock have a minor impact on fault stability.
Figure 21: Scenarios W and H: $\chi_{max}$ over time for all faults after $P$, shown for CO2 (upper row) and H2 (lower row). Scenario W corresponds to reservoir weakening ($E$ decreased by 30% after $P$), whereas Scenario H represents reservoir hardening ($E$ increased by 30% after $P$).
@@ 4.1 Comparison with previous studies
A few papers addressing the topics of our work have recently been published. Works by Wassing et al. (2017) and Haug et al. (2018) are of particular interest because they focus on Rotliegend reservoirs in the Netherlands and northern Germany, providing insights into fault reactivation and fault rupture during PP. Although the authors: i) do not consider induced seismicity during UGS stages, ii) use a simplified 2D setting, and iii) employ a different modeling approach concerning fault activation, these studies carry out a parametric analysis on the same geometric/geomechanical features investigated in this paper.
@@
The findings of our work mostly align with the outcomes by Wassing et al. (2017) regarding fault failure mechanisms and slip initiation. Our study confirms that fault slip initiates at the top of the reservoir rocks where the initial shear capacity utilization SCU (Wassing et al., 2017), corresponding to our safety factor $\chi$, reaches the maximum values. Both studies agree that faults without offset experience reactivation at later stages of depletion, and fault rupture does not extend upwards into the caprock. Scenarios #2c, #2d, and C2, for example, confirm that in the case of a compartment offset, an earlier fault activation is favored and remains bounded within the reservoir depth range. Regarding the influence of the contrast in elastic properties distribution, our study agrees with Wassing et al. (2017) only partially. They observed a secondary peak in shear stress at the bottom of the footwall reservoir block and a localized decrease in shear stresses and SCU at the bottom of the hanging wall reservoir block. Qualitatively, the results presented above (see, for example, Figure 20) agree. However, it must be considered that Wassing et al. (2017) assigned an uniform density and elastic properties to all rocks while our study, in particular scenarios #4a and #4b, demonstrates that the contrast in elastic properties may significantly impact on stress concentration and, therefore, fault failure (Figure 11).
@@
Comparing our results with the findings presented by Haug et al. (2018) regarding factors influencing fault reactivation and criticality, we agree that the initial stress regime (scenarios #6a and #6b, Figure  15) plays significant roles in fault behavior. Obviously, this relation is a well knows in the literature (Foulger et al., 2018; Muntendam-Bos et al., 2015; Walsh and Zoback, 2016). Indeed, fault criticality largely increases as the horizontal components of the natural stress regime decrease relative to the vertical stress. Furthermore, our outcomes confirm that the stiffness contrast between the reservoir and surrounding rocks (scenarios #4a and #4b) governs the stress redistribution and the degree of stress rotation during the reservoir development, impacting fault reactivation. According to Haug et al. (2018) the depletion of thicker reservoir horizons results in a stronger fault-loading compared to the depletion of thinner reservoir horizons, as a thick reservoir undergoes a relatively larger strain for a same change of pore fluid pressure. Although our sensitivity analysis did not directly consider reservoir thickness as a parameter, we observed consistent outcomes when a larger strain is attributed to a larger pressure decrease (scenario #5b, Figure 14).
@@
However, there are a few aspects where our project disagrees with findings from Haug et al. (2018). While they suggested that the fault throw should be half of the reservoir thickness to obtain the maximum shear stress ratio values, we found that the most critical condition occurred when the fault offset is equal to the entire reservoir thickness. This discrepancy can be attributed to the different modeling setup, including the horizontal-to-vertical ratio of the natural stress components and the reservoir depth. These disparities emphasize the importance of the modeling approach and setup parameters, highlighting the need for further research to fully understand fault behavior in different geological contexts.